OpenStudy (anonymous):

Solve 2(x-1)(x-8)=0 PlEaSe HELP? :)

OpenStudy (asnaseer):

if (something) times (something else) is zero, then one of those must be zero

jimthompson5910 (jim_thompson5910):

Hint: 2(x-1)(x-8)=0 x-1=0 or x-8=0 By using the zero product property

OpenStudy (asnaseer):

agreed?

OpenStudy (anonymous):

Agreed^

OpenStudy (asnaseer):

so, in your case, we have:\[2(x-1)(x-8)=0\]which means either: 2 = 0, or x - 1 = 0, or x - 8 = 0

OpenStudy (asnaseer):

now obviously 2 cannot equal zero, so we can eliminate that

OpenStudy (anonymous):

x=1,x=8?

OpenStudy (asnaseer):

which leave us with: x - 1 = 0, or x - 8 = 0

OpenStudy (asnaseer):

correct!

OpenStudy (anonymous):

Thanks!

OpenStudy (asnaseer):

yw :)